Machinery purchased for $61,800 by Swifty Co. in 2016 was originally estimated to have a life of 8 years with a salvage value of $4,120 at the end of that time. Depreciation has been entered for 5 years on this basis. In 2021, it is determined that the total estimated life should be 10 years with a salvage value of $4,635 at the end of that time. Assume straight-line depreciation.

Partially correct answer. Your answer is partially correct. Try again.
Prepare the entry to correct the prior years' depreciation, if necessary. (If no entry is required, select "No entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Account Titles and Explanation

Debit

Credit

enter an account title

Entry field with incorrect answer now contains modified data

enter a debit amount

Entry field with incorrect answer

enter a credit amount

Entry field with correct answer

enter an account title

Entry field with incorrect answer now contains modified data

enter a debit amount

Entry field with correct answer

enter a credit amount

Entry field with incorrect answer

SHOW LIST OF ACCOUNTS

LINK TO TEXT

LINK TO TEXT

LINK TO VIDEO

Incorrect answer. Your answer is incorrect. Try again.
Prepare the entry to record depreciation for 2021. (If no entry is required, select "No entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually. Use Machinery related account.)

Account Titles and Explanation

Debit

Credit

enter an account title

Entry field with incorrect answer

enter a debit amount

Entry field with incorrect answer

enter a credit amount

Entry field with incorrect answer

enter an account title

Entry field with incorrect answer

enter a debit amount

Entry field with incorrect answer

enter a credit amount

Entry field with incorrect answer
Answer #1

Annual depreciation = (61800-4120)/8 = 7210

Net book value at the wnd of 2020 = 61800 - 7210*5 = $25750

Depreciation for 2021 = (25750-4635)/5 = $4223

Account titles and explanation Debit Credit
No entry required
Depreciation expense 4223
Accumulated depreciation 4223